# Don’t Overthink it. Preferring Shorter Thinking Chains for Improved LLM Reasoning

### v1

@@ 4.1 The short-m@k method
The short-m@k method, visualized in Figure 1, performs parallel decoding of $k$ generations for a given question, halting computation across all generations as soon as the $m\leq k$ shortest thinking trajectories are completed. It then conducts majority voting among those shortest answers, resolving ties by selecting the answer with the shortest thinking chain. Given that thinking trajectories can be computationally intensive, terminating all generations once the $m$ shortest trajectories are completed not only saves computational resources but also significantly reduces wall time due to the parallel decoding approach, as shown in Section 4.3.
Below we focus on short-1@k and short-3@k, with short-1@k being the most efficient variant of short-m@k and short-3@k providing the best balance of performance and efficiency (see Section 4.3). Ablation studies on $m$ and other design choices are presented in Section 4.4.
@@ 4.2 Evaluation setup
We evaluate all methods under the same setup as described in Section 3.1. We report results using our method (short-m@k) with $m\in\{1,3\}$. We compare the proposed method to the standard majority voting (majority@$k$), arguably the most common method for aggregating multiple outputs (Wang et al., 2022), which was recently adapted for reasoning LLMs (Guo et al., 2025; Abdin et al., 2025; Wang et al., 2025). As an oracle, we consider pass@$k$ (Kulal et al., 2019; Chen et al., 2021), which measures the probability of including the correct solution within $k$ generated responses.
@@
Figure 3: Comparing different inference methods under controlled thinking compute. short-1@k is highly performant in low compute regimes. short-3@k dominates the curve compared to majority@$k$.
We benchmark the different methods with sample sizes of $k\in\{1,2,...,10\}$, assuming standard parallel decoding setup, i.e., all samples are generated in parallel. For the oracle (pass@$k$) approach, we use the unbiased estimator presented in Chen et al. (2021), with our $20$ generations per question ($n$$=$$20$). For the short-1@k method, we use the rank-score@$k$ metric (Hassid et al., 2024), where we sort the different generations according to thinking length. For majority@$k$ and short-m@k where $m>1$, we run over all $k$-sized subsets out of the $20$ generations per example.
We evaluate the different methods considering three main criteria: (a) Sample-size (i.e., $k$), where we compare methods while controlling for the number of generated samples; (b) Thinking-compute, where we measure the total number of thinking tokens used across all generations in the batch; and (c) Time-to-answer, which measures the wall time of running inference using each method. We note that we assume that currently served LLMs operate with a fixed processing speed in a fixed batch setting while using parallel decoding. As a result, using our method (short-m@k), we terminate all other generations after the first $m$ decoding thinking processes terminate. Thus, the overall thinking compute is the total number of thinking tokens for each of the $k$ generations at that point. Similarly, the overall time is that of the $m$’th shortest generation process. Conversely, for majority@$k$, the method’s design necessitates waiting for all generations to complete before proceeding. Hence, we consider the compute as the total amount of thinking tokens in all generations and run time according to the longest thinking chain. As for the oracle approach, we terminate all thinking trajectories once the shortest correct one is finished, and consider the compute and time accordingly.
4.3 Results
@@
We start by examining different methods across benchmarks for a fixed sample size $k$. Results aggregated across benchmarks are presented in Figure 2, and detailed results per benchmark can be seen at Appendix A. We observe that, generally, all methods improve with larger sample sizes, indicating that increased generations per question enhance performance. This trend is somewhat expected for the oracle (pass@$k$) and majority@$k$ methods but surprising for our method, as it means that even when a large amount of generations is used, the shorter thinking ones are more likely to be correct. The only exception is QwQ-$32$B (Figure 2(c)), which shows a small of decline when considering larger sample sizes with the short-1@k method.
When comparing short-1@k to majority@$k$, the former outperforms at smaller sample sizes, but is outperformed by the latter in two out of three models when the sample size increases. Meanwhile, the short-3@k method demonstrates superior performance, dominating across nearly all models and sample sizes. We next analyze how this performance advantage translates into efficiency benefits.
@@
The aggregated performance results, evaluated with respect to thinking compute, are presented in Figure 3 (per-benchmark results provided in Appendix A). We again observe that the short-1@k method outperforms majority@$k$ at lower compute budgets. Notably, for LN-Super-$49$B (Figure 3(a)), the short-1@k method surpasses majority@$k$ across all compute budgets. For instance, short-1@k achieves $57\%$ accuracy with approximately $60\%$ of the compute budget used by majority@$k$ to achieve the same accuracy. For R1111-$32$B and QwQ-$32$B models, the short-1@k method exceeds majority@$k$ up to compute budgets of $45$k and $60$k total thinking tokens, respectively, but is underperformed by it on larger compute budgets.
@@
The short-3@k method yields even greater performance improvements, incurring only a modest increase in thinking compute compared to short-1@k. When compared to majority@$k$, short-3@k consistently achieves higher performance with lower thinking compute across all models and compute budgets. For example, with the QwQ-$32$B model (Figure 3(c)), and an average compute budget of $80$k thinking tokens per example, short-3@k improves accuracy by $2\%$ over majority@$k$.
@@
Figure 4: Comparing time-to-answer for different inference methods. Our methods substantially reduce time cost with no major loss in performance. short-3@k improves performance across all models and sample sizes. We note that unlike majority@$k$, which becomes slower as $k$ grows, our methods run faster with $k$, as the probability of finding a short chain increases with $k$.
Time-to-answer.
Finally, the aggregated time-to-answer results are shown in Figure 4, with per-benchmark details in Appendix A. [^4] For readability, the oracle is excluded from Figure 4, and methods are compared across a subset of sample sizes. As sample size increases, majority@$k$ exhibits longer time-to-answer, driven by a higher probability of sampling generations with extended thinking chains, requiring all trajectories to complete. Conversely, the short-1@k method shows reduced time-to-answer with larger sample sizes, as the probability of encountering a short answer increases. This trend also holds for the short-3@k method after three reasoning processes complete.
This phenomenon makes the short-1@k and short-3@k methods substantially more usable compared to basic majority@$k$. For example, when using the LN-Super-$49$B model (Figure 4(a)), with a sample size of $5$, the short-1@k method reduces time consumption by almost $50\%$ while also increasing performance by about $1.5\%$ compared to majority@$k$. When considering a larger sample size of $9$, the performance values are almost the same but short-1@k is more than $55\%$ faster.
Finally, we observe that for most models and sample sizes, short-3@k boosts performance, while for larger ones it also reduces time-to-answer significantly. For example, on R1111-$32$B (Figure 4(b)), with $k=5$, short-3@k is 33% faster than majority@$k$, while reaching superior performance. A similar boost in time-to-answer and performance is observed with QwQ-$32$B and sample size $9$ (Figure 4(c)).
4.4 Ablation studies

### v2

@@ 4.1 The short-m@k method
The short-m@k method, visualized in Figure˜1, performs parallel decoding of $k$ generations for a given question, halting computation across all generations as soon as the $m\leq k$ shortest thinking trajectories are completed. It then conducts majority voting among those shortest answers, resolving ties by selecting the answer with the shortest thinking chain. Given that thinking trajectories can be computationally intensive, terminating all generations once the $m$ shortest trajectories are completed not only saves computational resources but also significantly reduces wall time due to the parallel decoding approach, as shown in Section˜4.3.
Below we focus on short-1@k and short-3@k, with short-1@k being the most efficient variant of short-m@k and short-3@k providing the best balance of performance and efficiency (see Section˜4.3). Ablation studies on $m$ and other design choices are presented in Appendix˜C, while results for smaller models are presented in Appendix˜D.
@@ 4.2 Evaluation setup
We evaluate all methods under the same setup as described in Section˜3.1. We report the averaged results across the math benchmarks, while the results for GPQA-D presented in Appendix˜A. The per benchmark resutls for the math benchmarks are in Appendix˜B. We report results using our method (short-m@k) with $m\in\{1,3\}$. We compare the proposed method to the standard majority voting (majority@k@k), arguably the most common method for aggregating multiple outputs (Wang et al., 2022), which was recently adapted for reasoning LLMs (Guo et al., 2025; Abdin et al., 2025; Wang et al., 2025b). As an oracle, we consider pass@k@k (Kulal et al., 2019; Chen and others, 2021), which measures the probability of including the correct solution within $k$ generated responses.
We benchmark the different methods with sample sizes of $k\in\{1,2,...,10\}$, assuming standard parallel decoding setup, i.e., all samples are generated in parallel. [^2] Section 5.3 presents sequential analysis where parallel decoding is not assumed. For the oracle (pass@$k$) approach, we use the unbiased estimator presented in Chen and others (2021), with our $20$ generations per question (nn==2020). For the short-1@k method, we use the rank-score@$k$ metric (Hassid et al., 2024), where we sort the different generations according to thinking length. For majority@k@k and short-m@k where $m>1$, we run over all $k$-sized subsets out of the $20$ generations per example.
We evaluate the different methods considering three main criteria: (a) Sample-size (i.e., $k$), where we compare methods while controlling for the number of generated samples; (b) Thinking-compute, where we measure the total number of thinking tokens used across all generations in the batch; and (c) Time-to-answer, which measures the wall time of running inference using each method. In this parallel framework, our method (short-m@k) terminates all other generations after the first $m$ decoding thinking processes terminate. Thus, the overall thinking compute is the total number of thinking tokens for each of the $k$ generations at that point. Similarly, the overall time is that of the $m$’th shortest generation process. Conversely, for majority@k@k, the method’s design necessitates waiting for all generations to complete before proceeding. Hence, we consider the compute as the total amount of thinking tokens in all generations and run time according to the longest thinking chain. As for the oracle approach, we terminate all thinking trajectories once the shortest correct one is finished, and consider the compute and time accordingly.
4.3 Results
@@
(d) R1-670B
Figure 3: Comparing different inference methods under controlled thinking compute. short-1@k is highly performant in low compute regimes. short-3@k dominates the curve compared to majority@k@k.
@@
We start by examining different methods across benchmarks for a fixed sample size $k$. Results aggregated across math benchmarks are presented in Figure˜2, while Figure˜6 in Appendix˜A presents GPQA-D results, and detailed results per benchmark can be seen at Appendix˜B. We observe that, generally, all methods improve with larger sample sizes, indicating that increased generations per question enhance performance. This trend is somewhat expected for the oracle (pass@$k$) and majority@$k$ methods but surprising for our method, as it means that even when a large amount of generations is used, the shorter thinking ones are more likely to be correct. The only exception is QwQ-3232B (Figure˜2(c)), which shows a small of decline when considering larger sample sizes with the short-1@k method.
When comparing short-1@k to majority@$k$, the former outperforms at smaller sample sizes, but is outperformed by the latter in three out of four models when the sample size increases. Meanwhile, the short-3@k method demonstrates superior performance, dominating across nearly all models and sample sizes. Notably, for the R11-670670B model, short-3@k exhibits performance nearly on par with the oracle across all sample sizes. We next analyze how this performance advantage translates into efficiency benefits.
@@
The aggregated performance results for math benchmarks, evaluated with respect to thinking compute, are presented in Figure˜3 (per-benchmark results provided in Appendix˜B), while the GPQA-D respective results are presented in Figure˜7 in Appendix˜A. We again observe that the short-1@k method outperforms majority@k@k at lower compute budgets. Notably, for LN-Super-4949B (Figure˜3(a)), the short-1@k method surpasses majority@k@k across all compute budgets. For instance, short-1@k achieves $57\%$ accuracy with approximately $60\%$ of the compute budget used by majority@$k$ to achieve the same accuracy. For R11-3232B, QwQ-3232B and R11-670670B models, the short-1@k method exceeds majority@k@k up to compute budgets of 4545k, 6060k and 100100k total thinking tokens, respectively, but is underperformed by it on larger compute budgets.
@@
The short-3@k method yields even greater performance improvements, incurring only a modest increase in thinking compute compared to short-1@k. When compared to majority@k@k, short-3@k consistently achieves higher performance with lower thinking compute across all models and compute budgets. For example, with the QwQ-3232B model (Figure˜3(c)), and an average compute budget of 8080k thinking tokens per example, short-3@k improves accuracy by $2\%$ over majority@$k$. For the R11-670670B model (Figure˜3(d)), short-3@k consistently outperforms majority voting, yielding an approximate $4\%$ improvement with an average token budget of 100100k.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 4: Comparing time-to-answer for different inference methods. Our methods substantially reduce time cost with no major loss in performance. Unlike majority@k@k, which becomes slower as $k$ grows, our methods run faster with $k$, as the probability of finding a short chain increases with $k$.
Time-to-answer.
Finally, the math aggregated time-to-answer results are shown in Figure˜4, with GPQA-D results shown in Figure˜8 and pe math benchmark in Appendix˜B. [^3] For readability, Figure 4 omits the oracle, and methods are compared across a subset of sample sizes. As sample size increases, majority@k@k exhibits longer time-to-answer, driven by a higher probability of sampling generations with extended thinking chains, requiring all trajectories to complete. Conversely, the short-1@k method shows reduced time-to-answer with larger sample sizes, as the probability of encountering a short answer increases. This trend also holds for the short-3@k method after three reasoning processes complete.
This phenomenon makes the short-1@k and short-3@k methods substantially more usable compared to basic majority@k@k. For example, when using the LN-Super-4949B model (Figure˜4(a)), with a sample size of $5$, the short-1@k method reduces time consumption by almost $50\%$ while also increasing performance by about $1.5\%$ compared to majority@$k$. When considering a larger sample size of $9$, the performance values are almost the same but short-1@k is more than $55\%$ faster.
Finally, we observe that for most models and sample sizes, short-3@k boosts performance, while for larger ones it also reduces time-to-answer significantly. For example, on R11-3232B (Figure˜4(b)), with $k=5$, short-3@k is $33\%$ faster than majority@$k$, while reaching superior performance. A similar boost in time-to-answer and performance is observed with QwQ-3232B/R11-670670B and sample size $9$ (Figure˜4(c) and Figure˜4(d)).
5 Analysis
To obtain deeper insights into the underlying process, making shorter thinking trajectories preferable, we conduct additional analysis. We first investigate the relation between using shorter thinking per individual example, and the necessity of longer trajectories to solve more complex problems (Section˜5.1). Subsequently, we analyze backtracks in thinking trajectories to better understand the characteristics of shorter trajectories (Section˜5.2). Lastly, we analyze the perfomance of short-m@k in sequential setting (Section˜5.3). All experiments in this section use trajectories produced by our models as described in Section˜3.1.444For Sections 5.1 and 5.2 we exclude generations that were not completed within the generation length.
5.1 Hard questions (still) require more thinking
We split the questions into three equal size groups according to model’s success rate. Then, we calculate the average thinking length for each of the splits, and provide the average lengths for the correct and incorrect attempts per split.
Table 2: Average thinking tokens for correct (C), incorrect (IC) and all (A) answers, per split by difficulty for the math benchmarks. The numbers are in thousands of tokens.
Model Easy Medium Hard
C/IC/A C/IC/A C/IC/A
LN-Super-49B $\phantom{0}5.3/11.1/\phantom{0}5.7$ $11.4/17.1/14.6$ $12.4/16.8/16.6$
R1-32B $\phantom{0}4.9/13.7/\phantom{0}5.3$ $10.9/17.3/13.3$ $14.4/15.8/15.7$
QwQ-32B [^5] The QwQ-32B and R1-670B models correctly answered all of their easier questions in all attempts. $\phantom{0}8.4/\phantom{0.}\text{--}\phantom{0}/\phantom{0}8.4$ $14.8/21.6/15.6$ $19.1/22.8/22.3$
R1-670Bfootnotemark: $13.0/\phantom{0.}\text{--}\phantom{0}/13.0$ $15.3/20.9/15.5$ $23.0/31.7/28.4$
Tables˜2 and 5 shows the averages thinking tokens per split for the math benchmarks and GPQA-D, respectively. We first note that as observed in Section˜3.2, within each question subset, correct answers are typically shorter than incorrect ones. This suggests that correct answers tend to be shorter, and it holds for easier questions as well as harder ones.
Nevertheless, we also observe that models use more tokens for more challenging questions, up to a factor of $2.9$. This finding is consistent with recent studies (Anthropic, 2025; OpenAI, 2024; Muennighoff et al., 2025) indicating that using longer thinking is needed in order to solve harder questions. To summarize, harder questions require a longer thinking process compared to easier ones, but within a single question (both easy and hard), shorter thinking is preferable.
5.2 Backtrack analysis
One may hypothesize that longer thinking reflect a more extensive and less efficient search path, characterized by a higher degree of backtracking and “rethinking”. In contrast, shorter trajectories indicate a more direct and efficient path, which often leads to a more accurate answer.
To this end, we track several keywords identified as indicators of re-thinking and backtracking within different trajectories. [^6] The keywords we used are: [’but’, ’wait’, ’however’, ’alternatively’, ’not sure’, ’going back’, ’backtrack’, ’trace back’, ’hmm’, ’hmmm’] We then categorize the trajectories into correct and incorrect sets, and measure the number of backtracks and their average length (quantified by keyword occurrences divided by total thinking length) for each set. We present the results for the math benchmarks and GPQA-D in Tables˜3 and 6, respectively.
Table 3: Average number of backtracks, and their average length for correct (C), incorrect (IC) and all (A) answers in math benchmarks.
Model # Backtracks Backtrack Len.
C/IC/A C/IC/A
LN-Super-49B $106/269/193$ $\phantom{0}88/\phantom{0}70/76$
R1-32B $\phantom{0}95/352/213$ $117/\phantom{0}63/80$
QwQ-32B $182/269/193$ $\phantom{0}70/\phantom{0}60/64$
R1-670B $188/323/217$ $\phantom{0}92/102/99$
As our results indicate, for all models and benchmarks, correct trajectories consistently exhibit fewer backtracks compared to incorrect ones. Moreover, in almost all cases, backtracks of correct answers are longer. This may suggest that correct solutions involve less backtracking where each backtrack is longer, potentially more in-depth that leads to improved reasoning, whereas incorrect ones explores more reasoning paths that are abandoned earlier (hence tend to be shorter).
Lastly, we analyze the backtrack behavior in a length-controlled manner. Specifically, we divide trajectories into bins based on their length. Within each bin, we compare the number of backtracks between correct and incorrect trajectories. Our hypothesis is that even for trajectories of comparable length, correct trajectories would exhibit fewer backtracks, indicating a more direct path to the answer. The results over the math benchmarks and GPQA-D are presented in Appendix˜F. As can be seen, in almost all cases, even among trajectories of comparable length, correct ones show a lower number of backtracks. The only exception is the R11-670670B model over the math benchmarks. This finding further suggests that correct trajectories are superior because they spend less time on searching for the correct answer and instead dive deeply into a smaller set of paths.
5.3 short-m@k with sequential compute
Our results so far assume sufficient resources for generating the output in parallel. We now study the potential of our proposed method without this constraint, by comparing short-m@k to the baselines in sequential (non-batched) setting, and measuring the amount of thinking tokens used by each method. For short-m@k, each generation is terminated once its length exceeds the maximum length observed among the $m$ shortest previously completed generations.
The results for the math benchmarks are presented in Figure˜5 while the GPQA-D results are in Appendix˜E. While the performance of short-m@k shows decreased efficiency in terms of total thinking compute usage compared to a fully batched decoding setup (Section˜4.3), the method’s superiority over standard majority voting remains. Specifically, at low compute regimes, both short-1@k and short-3@k demonstrate higher efficiency and improved performance compared to majority voting. As for higher compute regimes, short-3@k outperforms the majority voting baseline.
@@ 5.2 Backtrack analysis
(d) R1-670B
Figure 5: Comparing different methods for the math benchmarks under sequential (non-parallel) decoding.
6 Finetuning using shorter trajectories
Based on our findings, we investigate whether fine-tuning on shorter reasoning chains improves LLM reasoning accuracy. While one might intuitively expect this to be the case, given the insights from Section˜3 and Section˜5, this outcome is not trivial. A potential counterargument is that training on shorter trajectories could discourage the model from performing necessary backtracks (Section˜5.2), thereby hindering its ability to find a correct solution. Furthermore, the benefit of using shorter trajectories for bootstrapping reasoning remains an open question.
To do so, we follow the S11 paradigm, which fine-tunes an LLM to perform reasoning using only 1,1,$000$ trajectories (Muennighoff et al., 2025). We create three versions of the S11 dataset, built from examples with the shortest, longest, and random reasoning chains among several generations.
Data creation and finetuning setup.
To construct the three variants of S11, we generate multiple responses for each S11 question-answer pair. Specifically, for each example, we produce $10$ distinct answers using the QwQ-3232B model, which we select for its superior performance with respect to model size (Section˜3). From these $10$ responses per example, we derive three dataset variants—S11-short, S11-long, and S11-random—by selecting the shortest/longest/random response, respectively. This results in three datasets, each containing the same 1,1,$000$ queries but with distinct reasoning trajectories and answers. We then finetune the Qwen-2.52.5-77B-Instruct and Qwen-2.52.5-3232B-Instruct models on the three S11 variants. We further detail about the generation process, the finetuning setup and evaluation setup in Appendix˜G.
Finetuning results.
Results for the 32B model are presented in Table˜4 (7B model results are in Table˜10). For GPQA-D, AIME $2025$ and HMMT benchmarks, the S11-short variant achieves superior performance while using fewer thinking tokens. While performance on AIME $2024$ is similar across models, S11-short still demonstrates the shortest thinking. Aggregated results across math benchmarks reveal that S11-short improves relative performance by $2.8\%$ compared to the S11-random baseline, with a reduction of $5.8\%$ in thinking tokens. Conversely, the S11-long model consumes more tokens than S11-random, but obtains similar performance.
These results suggest that training on shorter reasoning sequences can lead to better reasoning models that exhibit reduced computational overhead. This observation aligns with our findings in Section˜3, which shows that answers with shorter thinking trajectories tend to be more accurate. We believe that developing models that reason more effectively with less computation holds significant potential.
Table 4: Results for our finetuned models over the S11 variants using Qwen-2.5-32B-Instruct: S11-short/long/random. The S11-short model improves performance over the other two models, while using fewer thinking tokens.
GPQA-D AIME 2024 AIME 2025 HMMT Math Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
S1-random 11566 62.5 16145 68.8 17798 59.3 19243 40.8 17729 56.3
[…]
7 Conclusion
In this work, we challenged the common assumption that increased test-time computation leads to better performance in reasoning LLMs. Through empirical analysis on four complex mathematical and reasoning benchmarks, we showed that shorter reasoning chains consistently outperform longer ones, both in accuracy and computational efficiency. Building on this insight, we introduced short-m@k, an inference method that prioritizes early-terminating generations. short-1@k, our most efficient variant, is preferred over traditional majority voting in low-compute settings. short-3@k, while slightly less efficient, outperforms majority voting across all compute budgets. We further investigate thinking trajectories, and find that shorter thinking usually involve less backtracks, and a more direct way to solution. To further validate our findings, we fine-tuned an LLM on short reasoning trajectories and observed improved accuracy and faster runtime, whereas training on longer chains yields diminishing returns. These findings highlight a promising direction for developing faster and more effective reasoning LLMs by embracing brevity over extended computation.
Acknowledgments
Appendix A GPQA diamond results
We present below results for the GPQA-D benchmark. Figure˜6, Figure˜7 and Figure˜8 are the sample-size/compute/time-to-answer results for GPQA-D, respectively. Table˜5 correspond to the Table˜2 in Section˜5.1. Table˜6 and Table˜9 correspond to Table˜3 and Table˜8 in Section˜5.2, respectively.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 6: GPQA-D - sample size ($k$) comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 7: GPQA-D - thinking compute comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 8: GPQA-D - time-to-answer comparison.
Table 5: Average thinking tokens for correct (C), incorrect (IC) and all (A) answers, per split by difficulty for GPQA-D. The numbers are in thousands of tokens.
Model Easy Medium Hard
C/IC/A C/IC/A C/IC/A
LN-Super-49B $2.5/\text{--}/2.5$ $\phantom{0}6.2/\phantom{0}7.8/\phantom{0}6.6$ $\phantom{0}7.1/\phantom{0}6.9/\phantom{0}7.0$
R1-32B $3.4/\text{--}/3.4$ $\phantom{0}6.4/\phantom{0}7.9/\phantom{0}6.8$ $\phantom{0}8.3/\phantom{0}7.8/\phantom{0}7.9$
QwQ-32B $5.3/\text{--}/5.3$ $\phantom{0}8.9/13.0/\phantom{0}9.7$ $11.1/10.6/10.6$
R1-670B $8.1/\text{--}/8.1$ $10.9/16.0/11.4$ $17.9/17.9/17.9$
Table 6: Average number of backtracks, and their average length for correct (C), incorrect (IC) and all (A) answers in GPQA-D.
Model # Backtracks Backtrack Len.
C/IC/A C/IC/A
LN-Super-49B $\phantom{0}89/107/\phantom{0}94$ $72/56/63$
R1-32B $\phantom{0}92/173/120$ $78/48/60$
QwQ-32B $152/241/178$ $52/41/46$
R1-670B $122/237/156$ $83/60/69$
Appendix B Per benchmark results
We present the per-benchmark results for each of the criteria presented in Section˜4.2. The sample-size ($k$) results are presented in Figures˜9, 10 and 11. The thinking compute comparison results are presented in Figures˜12, 13 and 14. The time-to-answer results per benchamrk are presented in Figures˜15, 16 and 17.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R11-670670B
Figure 9: AIME 2024 - sample size ($k$) comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 10: AIME 2025 - sample size ($k$) comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 11: HMMT Feb 2025 - sample size ($k$) comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 12: AIME 2024 - thinking compute comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 13: AIME 2025 - thinking compute comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 14: HMMT Feb 2025 - thinking compute comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 15: AIME 2024 - time-to-answer comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 16: AIME 2025 - time-to-answer comparison.
(a) LN-Super-4949B
(b) R11-3232B
(c) QwQ-32B
(d) R1-670B
Figure 17: HMMT Feb 2025 - time-to-answer comparison.
Appendix C Ablation studies
We investigate two axis of short-m@k: the value of $m$ and the tie breaking method. For all experiments we use LN-Super-4949B, reporting results over the three benchmarks described in Section˜3.1. For the ablation studies we focus on controlling thinking compute.
We start by ablating different $m\in\{1,3,4,5,7,9\}$ for short-m@k. Results are shown in Figure˜18(a). As observed in our main results, short-1@k outperforms others in low-compute regimes, while being less effective for larger compute budgets. Larger $m$ values seem to perform similarly, with higher $m$ values yielding slightly better results in high-compute scenarios.
Next, we analyze the tie-breaking rule of short-m@k. We suggest the selection of the shortest reasoning chain among the vote-leading options. We compare this strategy to random tie-breaking, and to tie breaking according to the longest reasoning chain among the options. As shown in Figure˜18(b), the short-m@k strategy outperforms random tie-breaking. In contrast, choosing the option with the longest reasoning chain yields inferior results.